OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

pooja195 (pooja195):

Divide both sides by -3. −3(|2x+6|)/-3=−12/−3 |2x+6|=4 Now just set up two equations (neg and pos solution) 2x+6=-4 2x+6=4 then just solve for x you will have 2 solutions